# modelscsfList of Model Score Tables {CJK*}UTF8gbsn

### v1

@@ Appendix C Multilingual Subjective Evaluation
Every road you take, I’ll be one step behind
@@
Every road you take, I’ll be one step behind
@@
I’ll cross the oceans wide, stand by your side
@@
Every road you take, I’ll be one step behind
@@
Promise me you’ll stay awhile, don’t ever say goodbye
Together we’ll chase every star across the sky
@@
Hoping someday you’ll feel it too
@@
That you’ll come back and choose to stay
@@
In the love we’ve made, we’ll find the light
@@
You teach me yours and I’ll show you mine
@@
Together we’ll rewrite our history
@@
We’ll be getting down tonight until the morning light
@@
Better love, we’ll be making
@@
We’ll keep the flame forever here.
@@
We’ll keep the flame forever here.
@@
We’ll keep the flame forever here.
@@
We’ll keep the flame forever here.
@@
Turn it up, we’ll sing aloud.
@@
Take my hand, we’ll show them how.
@@
Every road you take, I’ll be one step behind
@@
Every road you take, I’ll be one step behind
@@
I’ll cross the oceans wide, stand by your side
@@
Every road you take, I’ll be one step behind
@@
Promise me you’ll stay awhile, don’t ever say goodbye
Together we’ll chase every star across the sky
@@
Hoping someday you’ll feel it too
@@
That you’ll come back and choose to stay

### v2

@@ 14 Ethics and Responsibility
This study has been reviewed and approved by the Human and Artefacts Research Ethics Committee under protocol HREP-2023-0230, titled Building Platform Technologies for Symbiotic Creativity in Hong Kong. The approval ensures that our research adheres to ethical guidelines in data usage, AI generation, and cultural representation. The approval remains effective until 30-Jan-2027.
@@ Appendix C Multilingual Subjective Evaluation
Every road you take, I’$l$ be one step behind
@@
Every road you take, I’$l$ be one step behind
@@
I’$l$ cross the oceans wide, stand by your side
@@
Every road you take, I’$l$ be one step behind
@@
Promise me you’$l$ stay awhile, don’t ever say goodbye
Together we’$l$ chase every star across the sky
@@
Hoping someday you’$l$ feel it too
@@
That you’$l$ come back and choose to stay
@@
In the love we’ve made, we’$l$ find the light
@@
You teach me yours and I’$l$ show you mine
@@
Together we’$l$ rewrite our history
@@
We’$l$ be getting down tonight until the morning light
@@
Better love, we’$l$ be making
@@
We’$l$ keep the flame forever here.
@@
We’$l$ keep the flame forever here.
@@
We’$l$ keep the flame forever here.
@@
We’$l$ keep the flame forever here.
@@
Turn it up, we’$l$ sing aloud.
@@
Take my hand, we’$l$ show them how.
@@
Every road you take, I’$l$ be one step behind
@@
Every road you take, I’$l$ be one step behind
@@
I’$l$ cross the oceans wide, stand by your side
@@
Every road you take, I’$l$ be one step behind
@@
Promise me you’$l$ stay awhile, don’t ever say goodbye
Together we’$l$ chase every star across the sky
@@
Hoping someday you’$l$ feel it too
@@
That you’$l$ come back and choose to stay